OpenStudy (anonymous):

PLEASE HELP!!!!!!!! GIVING MEDALS!!! 4.189 = x/360 x 18.84

OpenStudy (anonymous):

4.189= x/(360*18.84) ??

OpenStudy (anonymous):

I have no idea but I have to go in 5 minutes.. Can you just gibe me the answer?? I would really appreciate it... There's another one after this

OpenStudy (anonymous):

if so... then x= (360*18.84*4.189) x=28411.4736

OpenStudy (anonymous):

Thanks so much!!!! <3 Can you help with just one more?!

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

What is the arc measure of a sector with area 38.485 ft2 and radius 7?

OpenStudy (anonymous):

If the radius is 7, the entire circle would have area pi*7^2 = 49*pi =~ 153.938 But we only have 38.485, so divide that by (49*pi) and we have: 38.485 / (49*pi) =~ 0.25, that's almost exactly 1/4 of a circle (90 degrees or pi/2 radians)

OpenStudy (anonymous):

erm.. answer please... I don't have timeeeee

OpenStudy (anonymous):

And Thanks again. It does mean a lot..

OpenStudy (anonymous):

10.99 about

OpenStudy (anonymous):

Thanks so much! Good Night!!!!

OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm.

OpenStudy (anonymous):

For that one it would be the 18411.4736 correct?

OpenStudy (anonymous):

28411.4736***

OpenStudy (anonymous):

4.189/[2 π 3] = x/360 6 π x = 360 *4.189

OpenStudy (anonymous):

80

OpenStudy (anonymous):

80?!

OpenStudy (anonymous):

yup arc measure

OpenStudy (anonymous):

youve got arc length and radius. pi= 3.14

OpenStudy (anonymous):

4.189 x 360/2(3.14)(3)

OpenStudy (anonymous):

just abut 80